# Particles before symmetry

### v4

@@ 3.2 Dynamical Mass Generation for the Higgs in the Linearised Theory
We assume $E$ admits at least one unit section and pick one, $e_{0}(x)$. Two different choices $e_{0},e^{\prime}_{0}$ will be isomorphic: they are structurally indistinguishable within (their copies of) $E$, since they have the same (unit) norm (and there is no more background structure on $E$).
Thus, in the usual terminology, any two such choices can be mapped to each other by an element of $\mathsf{Aut}(E)$. So, in order to parametrise the space of non-isomorphic unit sections, we can fix the same $e_{0}$ across configurations. In the jargon of \@BBOPcitep\@BAP\@BBN(Gomes, 2025b; Kabel et al., 2025; Gomes & Butterfield, 2023)\@BBCP, we take $e_{0}$ to define a ‘representational scheme’ across physical possibilities, so that each configuration $\varphi\in\Gamma(E)$ is parallel to the same $e_{0}$. This is this demonstration’s analogue of choosing ‘unitary gauge’, and it is the only place where symmetry could be mentioned. [^14] Indeed, representational schemes can be compared to gauge-fixings (cf. \@BBOPcite\@BAP\@BBN(Gomes, 2025b, Sec. 3.3)\@BBCP), as follows. Consider $\Gamma(E)$, and its sector $\Gamma_{0}(E)$. Let $\varphi,\varphi^{\prime}\in\Gamma_{0}(E)$. The group $\mathsf{Aut}(E)$ acts transitively on the unit normal sections: it can take any internal direction into any other. Therefore, we could, by a suitable gauge transformation on $\varphi$, make it collinear with $\varphi^{\prime}{}$. Once they are collinear, it is a trivial matter to separate out the part that has a given norm, be it $e_{0}$, $v^{\prime}e_{0}$, or $ve_{0}$ (see below, for the definition of $v$).
@@
We thus would now like to see $H$ as parametrising the tangent vectors. But instead of thinking in terms of tangent bundles to configuration spaces, it is more practical to use the standard assumptions of ‘low energy’, or, equivalently, going into the linearised regime and using expansions in terms of order parameters, i.e. $\mathcal{O}(H)=\mathcal{O}(\nabla)=\mathcal{O}(K_{e_{0}})=\varepsilon$. [^15] In the comparative sense: that $\frac{|H|}{v}\sim\varepsilon<<1$, and mutatis mutandis for the appropriate norm on $\nabla$. The pointwise norm on the space of covariant derivatives can be defined in two equivalent ways: either by seeing it as an affine space over the space of connections $\omega\in\Gamma(T^{*}M\otimes\mathsf{End}(E))$, which requires us to pick a flat covariant derivative $\nabla_{o}$ and then take the norm of $\omega=\nabla-\nabla_{o}$; or, alternatively, for $v\in T_{x}M$ of unit $g$-norm, and $s\in E_{x}$ of unit $\langle\cdot,\cdot\rangle$-norm, one can define $\|\nabla\|_{\mathrm{op}}:=\sup_{v,s}\|\nabla_{v}s\|.$ Using $\varphi=(H+v)e_{0}$ to rewrite the kinetic term (3.5) we once again get
@@
The kinetic term for the Higgs is given by $(\partial H)^{2}$ whereas $v^{2}$ is the mass term for part of the affine structure. Now let us turn to the mass of the Higgs, for which we need the Higgs potential.
@@
where here $V:\Gamma(E)\rightarrow\mathbb{R}$ is assumed to be homogeneous on $M$ and depend only on $\|\varphi\|$, and so $w_{o}(x)=v\,e_{0}(x)\in\Gamma_{0}(E)$ with $\|w_{o}(x)\|=v\neq 0,\,\forall x\in M$.
In order to get the mass of the Higgs we must similarly expand $V(\varphi)$ up to $\mathcal{O}(\varepsilon)^{2}$, thus involving the second variation of $V$ around its minimum, called the Hessian. The Hessian is a symmetric linear map that acts fiber-wise; using the notation $\delta\varphi$ for elements of $T_{w_{o}}\Gamma_{0}(E)$ we write it as:
@@
The mass terms will be obtained from (3.16) by diagonalising $\mathsf{Hess}(V)_{w_{o}}$. Thus, as in the standard account, the mass terms for different components of the Higgs along the different directions within $T_{w_{0}}\Gamma_{0}(E)$ are given by the eigenvalues of the Hessian with those directions as eigenvectors. But, per (3.10), in the current case we are considering only $\delta\varphi=H(x)e_{0}\in T_{w_{o}}\Gamma_{0}(E)$: only $e_{0}$ can have non-vanishing eigenvectors (the eigenvalues of any completion of $e_{0}$ into an orthogonal basis for $\mathsf{Re}\,\langle\,.\,$$\rangle$ (see footnote 12) would vanish for $e_{I},I\neq 0$). The value of the surviving eigenvalue along $e_{0}$ will depend on the form of $V(\|\varphi\|)$, which is assumed to be translationally invariant (or homogeneous on $M$). Here we will just call this eigenvalue $m_{H}^{2}$. [^16] In the standard case where $V=-\mu\|\varphi\|^{2}+\lambda\|\varphi\|^{4}$, using (3.9) and (3.10) we can replace $\|\varphi\|^{2}\rightarrow(H+v)^{2}$, and $\frac{\delta}{\delta\varphi}\rightarrow\frac{\partial}{\partial H}$. So $\mathsf{Hess}(V)\rightarrow\frac{\partial^{2}V}{\partial H^{2}}$. Then an easy computation reveals: $\frac{\partial V}{\partial H}=-2\mu(H+v)+4\lambda(H+v)^{3}$ (3.17) and $\frac{\partial^{2}V}{\partial H^{2}}=-2\mu+12\lambda(H+v)^{2},$ (3.18) Setting $\frac{\partial V}{\partial H}=0$ and $H=0$ in (3.17) we obtain $v=\sqrt{\frac{\mu}{2\lambda}}$. Setting $v=\sqrt{\frac{\mu}{2\lambda}}$ and $H=0$ in (3.18) we obtain $\frac{\partial^{2}V}{\partial H^{2}}(w_{o})=4\mu=m_{H}^{2}.$
@@
This concludes the geometric gloss on the Higgs mechanism. No explicit mention of stabilisers, decomposition of gauge orbits, Killing forms on the Lie algebra, Goldstone’s theorem, etc was required—devices indispensable in the standard formulation (see \@BBOPcite\@BAP\@BBN(Hamilton, 2017, Ch. 8.1)\@BBCP; \@BBOPcite\@BAP\@BBN(Tong, 2025, Ch. 2.2)\@BBCP; \@BBOPcite\@BAP\@BBN(Bleecker, 1981, Ch. 10.3)\@BBCP for comparison). [^17] For instance, the disappearance of Goldstone modes through a choice of gauge is replaced here by having all $\varphi\in\Gamma_{0}(E)$ parallel to $e_{0}$ and the orthogonality relation (3.4).
@@ 4 The Yukawa mechanism
Whereas the Higgs mechanism is used to ‘endow mass’ to the gauge potentials, the Yukawa mechanism is used to endow mass to the matter fields.
In the Standard Model fermion masses cannot be introduced as they can for real or complex-valued scalar fields. First of all, a Dirac mass term must couple left- and right-handed chiral fermions; moreover, the two chiralities are mapped into internal spaces that transform differently under the gauge group $G=SU(3)\times SU(2)\times U(1)$, so coupling them would violate gauge invariance: this is related to the issue we saw in Section 2.1 about canonical isomorphisms between associated vector bundles with different representations. The solution is to introduce the Higgs field $\varphi$, in such a way that gauge invariance is preserved, while the fermions acquire effective masses. This is the Yukawa mechanism.
Here I will essentially follow the treatment given in \@BBOPcite\@BAP\@BBN(Hamilton, 2017, Ch. 8)\@BBCP, whose notation and general approach is already much closer to the geometric approach that I’m pursuing here (as compared to the treatment of more familiar textbooks, for instance, the one given in \@BBOPcitep\@BAP\@BBN(Tong, 2025; Skinner, 2007)\@BBCP, which use representation theory more heavily). So I will call the treatment to be followed here ‘the standard’ treatment of the Yukawa mechanism. In Section 4.1 I will describe the obstruction to the formulation of mass terms for fermions, and its resolution in this, geometric-friendly but still ‘standard’, exposition. Then in Section 4.2 I will discuss what I think is explanatorily unsatisfactory about this resolution, and say why I take the VB-POV to provide a more transparent explanation.
4.1 The ‘standard’ presentation of the Yukawa mechanism
In more detail, here is the obstruction to the formulation of mass terms for fermions. Fermions are spinors, but for Weyl spinors, the inner product is anti-diagonal in the left and right basis: $\overline{\psi}_{R}\psi_{R}=0$, and so, in order to extract mass terms we must couple left to right-handed spinors: $\overline{\psi}_{R}\psi_{L}$. Thus, if both $\psi_{L}$ and $\psi_{R}$ are valued in the same internal space, i.e. in the same vector bundle, and are in the same representation, one may add mass terms of the form:
@@ 4.1 The ‘standard’ presentation of the Yukawa mechanism
and this will be gauge invariant since $\psi_{L}$ and $\psi_{R}$ transform in the same representation of the gauge group. I.e. locally, $\psi_{L}\in\Gamma(S_{L}\otimes E),\,\psi_{R}\in\Gamma(S_{R}\otimes E)$, where $(E,M,V)$ is the vector bundle with the representation space $V$ of the gauge group in question, and $S_{L}$ is the bundle of left-handed spinors over spacetime, whose typical fibre space is called $\Delta_{L}$ (mutatis mutandis for right-handed spinors).
@@
antilinear in $V_{L}$, real linear in $W$, linear in $V_{R}$. What do these maps look like, more precisely? Let us look at an example. Consider the $SU(2)\times U(1)$ representations for the leptons (taken from Table 1):
@@
The map $\tau$ is defined on vector spaces, and therefore depends on the choice of trivialisation of the vector bundles, i.e. on local decompositions of the form $E_{|U}\simeq U\times V$. To render this construction invariant, we must extend $\tau$ to sections of the associated vector bundles. This is straightforward. Given a section $\sigma(x)$ of an $SU(2)\times U(1)$ principal bundle (cf. (2.3)), we can use the local maps $l_{L}:U\rightarrow V_{L}$, $\varphi:U\rightarrow W$, and $l_{R}:U\rightarrow V_{R}$ introduced above to form the corresponding global sections $e_{L}\in\Gamma(S_{L}\!\otimes\!E_{L})$, $\varphi\in\Gamma(F)$, and $e_{R}\in\Gamma(S_{R}\!\otimes\!E_{R})$, which are independent of any choice of trivialisation. For instance, a left-handed electron (I will discuss its ‘up’ and ‘down’ components shortly) is represented as
@@
There are, however, other possible invariant maps that achieve the same result, and it is not immediately clear how to choose among them. For instance, let $H$ denote the centraliser of GG—the subgroup of elements commuting with all $g\in G$. Then for any such map producing a scalar singlet, such as (4.8), one can define a family
@@
with $h\in H$. For $G=SU(2)\times U(1)$, the centraliser is simply $\mathbb{C}$, acting as multiples of the identity. Thus, the ambiguity here amounts to an arbitrary complex factor, which in practice could be absorbed into the Yukawa couplings. For more complicated groups, however, such ambiguities may not be so easily removed. [^18] For a single Higgs doublet with SM hypercharges, the residual $C^{\times}$ freedom in the centraliser can be absorbed by field rephasings and a Yukawa reparametrisation, hence is physically inessential. In multi-Higgs or extended-gauge settings, the centraliser may act non-trivially on distinct invariant maps, leaving physically inequivalent couplings. The crude impression is that in the PFB-POV one manufactures invariants in a chosen basis while absorbing ambiguities into coupling constants.
Eliminating such extraneous factors is a matter of conceptual housekeeping: the geometry-first perspective accomplishes this by treating the Yukawa couplings not as equivariant maps on the underlying representation spaces, but as natural operations between the relevant vector bundles, e.g. as given in Equation (2.2) in Section 2.8.
4.2 The VB-POV presentation of the Yukawa mechanism
In Section 2.1 I argued that there was no canonical relation between associated vector bundles corresponding to different representations of the principal bundle; but of course we don’t need such a map to get a scalar out of sections of these bundles. All we need is that $T$, given in (4.8), is a gauge-invariant map between associated vector bundles, with $\tau$ a map between the representation spaces; and presenting one such map is sufficient for comparison with experiments. Nonetheless, lacking the canonical relation, I find this answer unsatisfactory, because opaque: why this particular map? Couldn’t we have found others? And how should we interpret them?
I take the geometric, VB-POV, to provide a more transparent interpretation of what the map $T$ represents, and what other choices would represent. Again, in the geometry-first formulation, all we have are structures in the fundamental vector bundle spaces. The fundamental vector spaces are given by $(E^{n},M,\mathbb{C}^{n},\langle\cdot,\cdot\rangle_{n})$, for $n=1,2,3$ (we will include orientation as further structure below, when we look at the Yukawa form for quarks). Different particles are merely different sections of different tensor products for these fundamental vector spaces. We replace ‘quantum numbers’ by a geometric characterisation of a given particle. Thus, for instance, a down-right-handed quark (of any of the three generations, but here we assume the first) is given by:

### v5

@@ 3.2 Dynamical Mass Generation for the Higgs in the Linearised Theory
We assume $E$ admits at least one unit section and pick one, $e_{0}(x)$. Again, two different choices $e_{0},e^{\prime}_{0}$ will be isomorphic: they are structurally indistinguishable within (their copies of) $E$, since they have the same (unit) norm (and there is no more background structure on $E$). Thus, in the usual terminology, any two such choices can be mapped to each other by an element of $\mathsf{Aut}(E)$. So, in order to parametrise the space of non-isomorphic unit sections, we can fix the same $e_{0}$ across configurations. In the jargon of \@BBOPcitep\@BAP\@BBN(Gomes, 2025a; Kabel et al., 2025; Gomes and Butterfield, 2023)\@BBCP, we take $e_{0}$ to define a ‘representational scheme’ across physical possibilities, so that each configuration $\varphi\in\Gamma(E)$ is parallel to the same $e_{0}$. This is this demonstration’s analogue of choosing ‘unitary gauge’, and it is the only place where symmetry could be mentioned. [^16] Indeed, representational schemes can be compared to gauge-fixings (cf. \@BBOPcite\@BAP\@BBN(Gomes, 2025a, Sec. 3.3)\@BBCP), as follows. Consider $\Gamma(E)$, and its sector $\Gamma_{0}(E)$. Let $\varphi,\varphi^{\prime}\in\Gamma_{0}(E)$. The group $\mathsf{Aut}(E)$ acts transitively on the unit normal sections: it can take any internal direction into any other. Therefore, we could, by a suitable gauge transformation on $\varphi$, make it collinear with $\varphi^{\prime}{}$. Once they are collinear, it is a trivial matter to separate out the part that has a given norm, be it $e_{0}$, $v^{\prime}e_{0}$, or $ve_{0}$ (see below, for the definition of $v$).
@@
We thus would now like to see $H$ as parametrising the tangent directions to a certain functional space characterising the physically distinct configurations of the Higgs field. But instead of thinking in terms of tangent bundles to configuration spaces, it is more practical to use the standard assumptions of ‘low energy’, or, equivalently, going into the linearised regime and using expansions in terms of order parameters, i.e. $\mathcal{O}(H)=\mathcal{O}(\nabla)=\mathcal{O}(K_{e_{0}})=\varepsilon$. [^17] In the comparative sense: that $\frac{|H|}{v}\sim\varepsilon<<1$, and mutatis mutandis for the appropriate norm on $\nabla$. The pointwise norm on the space of covariant derivatives can be defined in two equivalent ways: either by seeing it as an affine space over the space of connections $\omega\in\Gamma(T^{*}M\otimes\mathsf{End}(E))$, which requires us to pick a flat covariant derivative $\nabla_{o}$ and then take the norm of $\omega=\nabla-\nabla_{o}$; or, alternatively, for $v\in T_{x}M$ of unit $g$-norm, and $s\in E_{x}$ of unit $\langle\cdot,\cdot\rangle$-norm, one can define $\|\nabla\|_{\mathrm{op}}:=\sup_{v,s}\|\nabla_{v}s\|.$ Using $\varphi=(H+v)e_{0}$ to rewrite the kinetic term (3.5) we once again get
@@
The kinetic term for the Higgs is given by $(\partial H)^{2}$ whereas $v^{2}$ is the mass term for part of the affine structure, as we saw. Now let us turn to the mass of the Higgs, for which we need the Higgs potential.
@@
where here $V:\Gamma(E)\rightarrow\mathbb{R}$ is assumed to be homogeneous on $M$ and depend only on $\|\varphi\|$, and so $w_{o}(x)=v\,e_{0}(x)\in\Gamma_{0}(E)$ with $\|w_{o}(x)\|=v\neq 0,\,\forall x\in M$. So any potential whose minimum is non-zero suffices for identifying $w_{o}(x)=v\,e_{0}(x)$ for some $v>0.$ The familiar ’Mexi  [^18] Usually, one takes the Higgs vacuum vector $w_{o}$ to break the symmetry, and then one expands the Higgs fields around this vacuum. But then one must find a way to get rid of the ‘Goldstone modes’: directions of the Higgs field that are parallel to the gauge orbit of the vacuum. This is where one invokes unitary gauge. But this is awkward: angles are preserved by gauge transformations, so how could a gauge transformation act to eliminate modes that are defined by their inner product with respect to the Higgs vacuum, which is assumed to be a physical object?
In order to get the mass of the Higgs we must similarly expand $V(\varphi)$ up to $\mathcal{O}(\varepsilon)^{2}$, thus involving the second variation of $V$ around its minimum, called the Hessian. The Hessian is a symmetric linear map that acts fibre-wise; using the notation $\delta\varphi$ for elements of $T_{w_{o}}\Gamma_{0}(E)$ we write it as:
@@
The mass terms will be obtained from (3.16) by diagonalising $\mathsf{Hess}(V)_{w_{o}}$. Thus, as in the standard account, the mass terms for different components of the Higgs along the different directions within $T_{w_{0}}\Gamma_{0}(E)$ are given by the eigenvalues of the Hessian with those directions as eigenvectors. But, per (3.10), in the current case we are considering only $\delta\varphi=H(x)e_{0}\in T_{w_{o}}\Gamma_{0}(E)$: only $e_{0}$ can have non-vanishing eigenvectors (the eigenvalues of any completion of $e_{0}$ into an orthogonal basis for $\mathsf{Re}\,\langle\,.\,$$\rangle$ (see footnote 14) would vanish for $e_{I},I\neq 0$). The value of the surviving eigenvalue along $e_{0}$ will depend on the form of $V(\|\varphi\|)$, which is assumed to be translationally invariant (or homogeneous on $M$). Here we will just call this eigenvalue $m_{H}^{2}$. [^19] In the standard case where $V=-\mu\|\varphi\|^{2}+\lambda\|\varphi\|^{4}$, using (3.9) and (3.10) we can replace $\|\varphi\|^{2}\rightarrow(H+v)^{2}$, and $\frac{\delta}{\delta\varphi}\rightarrow\frac{\partial}{\partial H}$. So $\mathsf{Hess}(V)\rightarrow\frac{\partial^{2}V}{\partial H^{2}}$. Then an easy computation reveals: $\frac{\partial V}{\partial H}=-2\mu(H+v)+4\lambda(H+v)^{3}$ (3.17) and $\frac{\partial^{2}V}{\partial H^{2}}=-2\mu+12\lambda(H+v)^{2},$ (3.18) Setting $\frac{\partial V}{\partial H}=0$ and $H=0$ in (3.17) we obtain $v=\sqrt{\frac{\mu}{2\lambda}}$. Setting $v=\sqrt{\frac{\mu}{2\lambda}}$ and $H=0$ in (3.18) we obtain $\frac{\partial^{2}V}{\partial H^{2}}(w_{o})=4\mu=m_{H}^{2}.$
@@
This concludes the geometric gloss on the Higgs mechanism. No explicit mention of stabilisers, decomposition of gauge orbits, Killing forms on the Lie algebra, Goldstone’s theorem, etc was required—devices indispensable in the standard formulation (see \@BBOPcite\@BAP\@BBN(Hamilton, 2017, Ch. 8.1)\@BBCP; \@BBOPcite\@BAP\@BBN(Tong, 2025, Ch. 2.2)\@BBCP; \@BBOPcite\@BAP\@BBN(Bleecker, 1981, Ch. 10.3)\@BBCP for comparison). [^20] For instance, the disappearance of Goldstone modes through a choice of gauge is replaced here by having all $\varphi\in\Gamma_{0}(E)$ parallel to $e_{0}$ and the orthogonality relation (3.4). This is just what you would expect from the geometry alone.
@@ 4 The Yukawa mechanism
The Higgs mechanism endows mass to gauge potentials; the Yukawa mechanism endows mass to matter fields.
In the Standard Model, fermion masses cannot be introduced as for scalar fields. A Dirac mass term must couple left- and right-handed chiral fermions, but these transform in inequivalent representations of $G=SU(3)\times SU(2)\times U(1)$. Coupling them directly would violate gauge invariance—the same obstruction we encountered in Section 2.1 concerning canonical maps between associated bundles. The solution is to introduce the Higgs field in such a way that gauge invariance is preserved while fermions acquire effective masses.
I follow \@BBOPcite\@BAP\@BBN(Hamilton, 2017, Ch. 8)\@BBCP, whose notation is already geometric. Section 4.1 presents the standard treatment; Section 4.2 argues that the geometry-first picture does better.
4.1 The standard presentation
Fermions are spinors, and for Weyl spinors the inner product is anti-diagonal in chirality: $\overline{\psi}_{R}\psi_{R}=0$. Mass terms must therefore couple left to right: $\overline{\psi}_{R}\psi_{L}$. If both chiralities inhabit the same vector bundle and transform in the same representation, one may write
@@ 4.1 The standard presentation
and this is gauge-invariant. Locally, $\psi_{L}\in\Gamma(S_{L}\otimes E)$ and $\psi_{R}\in\Gamma(S_{R}\otimes E)$, where $(E,M,V)$ is the vector bundle with representation space $V$, and $S_{L}$ is the bundle of left-handed spacetime spinors with typical fibre $\Delta_{L}$ (mutatis mutandis for right-handed).
@@
antilinear in $V_{L}$, real linear in $W$, linear in $V_{R}$.
For the leptons, the relevant $SU(2)\times U(1)$ representations are (from Table 1):
@@
The map $\tau$ is defined on vector spaces and depends on the choice of trivialisation. To render it invariant, we extend to sections of the associated bundles. Given a section $\sigma(x)$ of an $SU(2)\times U(1)$ principal bundle (cf. (2.3)), the local maps $l_{L},\varphi,l_{R}$ define global sections $e_{L}\in\Gamma(S_{L}\!\otimes\!E_{L})$, $\varphi\in\Gamma(F)$, and $e_{R}\in\Gamma(S_{R}\!\otimes\!E_{R})$, independent of trivialisation. A left-handed electron is
@@
There are, however, other invariant maps that also produce a gauge-invariant scalar, and it is not immediately clear how to choose among them. For instance, let $H$ denote the centraliser of GG—the subgroup of elements commuting with all $g\in G$. Then for any such map producing a scalar singlet, such as (4.8), one can define a family
@@
with $h\in H$. For $G=SU(2)\times U(1)$, the centraliser is simply $\mathbb{C}$, acting as multiples of the identity. Thus, the ambiguity here amounts to an arbitrary complex factor, which in practice could be absorbed into the Yukawa couplings. For more complicated groups, however, such ambiguities may not be so easily removed. The crude impression is that in the PFB-POV one manufactures invariants in a chosen basis while absorbing ambiguities into coupling constants.
4.2 The geometry-first presentation
In Section 2.1 I argued that there is no canonical map between associated bundles with different representations. Of course we do not need such a map to extract scalars—$T$ in (4.8) suffices for comparison with experiment. But the answer is unsatisfying: why this particular map? Could we have found others?
The geometry-first perspective eliminates such ambiguities by treating Yukawa couplings not as equivariant maps on representation spaces, but as natural operations between vector bundles—inner products and contractions of the kind given in Equation (2.8). All we have are structures on the fundamental bundles $(E^{n},M,\mathbb{C}^{n},\langle\cdot,\cdot\rangle_{n})$ for $n=1,2,3$; in particular, $E^{1}$ is fixed so that $E^{1\otimes 3}$ has hypercharge +1+1 (so $E^{1}$ itself corresponds to $\tfrac{1}{3}$). Different particles are sections of different tensor products; quantum numbers become geometric labels. A down-right-handed quark is